# From Pixels to Insights: A Survey on Automatic Chart Understanding in the Era of Large Foundation Models

### v2

@@ 4 Modeling Strategies
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, line width=0.9pt, draw=connect-line, rounded corners,align=left, minimum width=2em, s sep=5pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=6em,font=, where level=3font=, where level=4font=, where level=5font=, [ Chart Understanding Approaches, black,rotate=90,anchor=north,edge=pure [ Classification-based Methods (§4.1) , pure-middle, edge=pure, text width=11.7em [ Fixed Output Vocab, pure-middle, edge=pure, text width=10.3em [ IMG+QUES [1], Relation Networks [67], V-MODEQA [45] , pure-leaf, text width=22em, edge=pure ] ] [ Dynamic Encoding, pure-middle, edge=pure, text width=10.3em [ SANDY [1], PreFIL [68], LEAF-Net [43], PlotQA-Pipe [4] , pure-leaf, text width=22em, edge=pure ] ] [ Pre-trained, pure-middle, edge=pure, text width=10.3em [ STL-CQA [44], TasPas [69], VisionTaPas [3] , pure-leaf, text width=22em, edge=pure ] ] ] [ Generation-based Methods (§4.2), node-middle, edge=node, text width=11.7em [ Without Pre-training, node-middle, edge=node, text width=10.3em [ SciCap-Net [6], FigCap-Net [51], Chart2Text [52] , node-leaf, text width=22em, edge=node ] ] [ Pre-trained, node-middle, edge=node, text width=10.3em [ ChartT5 [70], Donut [26], Pix2Struct [71], MatCha [72],
UniChart [73], FigCaps-HF [56], ChartVE [14] , node-leaf, text width=22em, edge=node ] ] [ Large Vision-language Models, node-middle, edge=node, text width=10.3em [ LLaVA [17], mPLUG-Owl [74], mPLUG-DocOwl [75],
SPHINX [76], MMCA [48], ChartLlama [38], ChartAssistant [39],
ChartInstruct [77], Gemini [78], GPT-4V [16], Claude 3 [79] , node-leaf, text width=22em, edge=node ] ] ] [ Tool Augmentation (§4.3), graph-middle, edge=graph, text width=11.7em [ OCR-based, graph-middle, edge=graph, text width=10.3em [ CRAFT [80], Parseq [81], RARE [82], FAN [83], ABINet++ [84],
CRNN [85], Star-Net [86], GRCNN [87], Rosetta [88] ,graph-leaf, text width= 22em, edge=graph ] ] [ Non-OCR-based, graph-middle, edge=graph, text width=10.3em [ DePlot [10], Chart-to-Table [14], DOMINO [89], PromptChart [11] ,graph-leaf, text width= 22em, edge=graph ] ] ] ]
@@ 4.1 Classification-based Models
Classification-based methods have primarily been developed for the task of chart question answering, leveraging a combination of visual and textual features to understand and interpret different types of charts. This subsection provides an overview of the evolution and key strategies employed in classification-based chart understanding.
@@ 4.2.2 Pre-trained Models
Significantly advancing the field, OCR-free end-to-end visual document understanding models like Donut [26] and Pix2Struct [71] introduced the capability to directly parse textual information from images, removing the need for separate OCR processes. Donut utilizes a Swin Transformer [98] for the image encoding and the decoder of mBART [99] for decoding, focusing on natural image parsing to produce text sequences from images. Pix2Struct, employing a Vision Transformer (ViT) [100] as the image encoder and a Transformer-based decoder. The pre-training objective is screenshot parsing, which learns to predict an HTML-based parse from a screenshot of a webpage, providing a structured training signal that encapsulates both textual and layout information. There are two major distinctions between Donut and Pix2Struct: (1) screenshot parsing may allow the model to learn the structure of the image better since HTML provides stronger signals about the layouts of the input image. (2) Donut is multilingual since it uses mBART’s decoder for decoding texts.
@@
First, do we still need external OCR systems? The advancements in chart understanding and interpretation achieved by UniChart and MatCha over ChartT5 highlight a pivotal shift in the field towards leveraging OCR-free models such as Pix2Struct and Donut. These developments underscore a crucial insight: the integration of advanced OCR-free pre-training objectives can obviate the dependency on traditional OCR systems, which often represent a bottleneck in terms of accuracy and adaptability. UniChart and MatCha’s superior performance, even without the direct input of ground truth tables that ChartT5 relies upon, evidently demonstrates the increased efficacy and robustness of models developed on the OCR-free paradigm.
@@
TABLE IV: Summary of various pre-trained and instruction fine-tuned vision-language models. Note that MMCA is instruction-tuned from mPLUG-Owl, ChartLlama is instruction-tuned from LLaVA-1.5, and ChartAssistant-S is instruction-tuned from SPHINX. The weights of ChartInstruct’s vision encoder are initialized from UniChart.
@@ 4.2.3 Large Vision-language Models
While the aforementioned pre-trained approaches significantly enhance chart understanding, they require task-specific fine-tuning (see Figure 3), limiting their generalizability. Contrastingly, large foundation models have demonstrated strong zero-shot or few-shot capabilities. For instance, in text generation, large language models (LLMs) like GPT-3 [42] showcase remarkable generalization to novel tasks when the model size and pre-training data reach certain scales. These models are able to handle new tasks in a zero-shot or few-shot manner, obviating the need for model retraining. Extending this paradigm into image understanding, studies have stacked vision encoders on top of LLMs, including methods such as LLaVA [17] and mPLUG-Owl [74], significantly advancing this field. Unique to these larger vision-language models (LVLMs), in contrast to the smaller vision-language models introduced previously, is the inclusion of a visual-language projector, $\textsc{Projector}_{\text{vt}}$, given the vision encoders and LLMs’ pre-training scales. Here, the visual representations obtained from the vision encoder are fed to $\textsc{Projector}_{\text{vt}}$ prior to concatenated with text representations.
@@
Finally, is it possible for models trained exclusively on synthetically generated data to match or exceed the performance of those trained on datasets curated from human-generated content? While synthetic data offers the advantage of scalability and control over diversity and complexity, it may lack the nuanced and unpredictable variations present in human-generated charts. Exploring whether synthetic data can truly capture the breadth of human creativity and intricacy in chart design is essential, as is understanding the potential gaps that might arise when such models are applied to real-world data.
@@ 5.2 Evaluation
Fairness and Bias. Evaluating fairness and mitigating biases in the generation of models’ content is critical when deploying these tools in user-facing applications. There are two main forms of bias to be aware of: fairness bias, which concerns the equitable representation and treatment of various groups (e.g., race, culture, or gender) [28, 119], and sample bias, which occurs when the data used to train the model is not representative of the broader population or scenario it’s intended to serve (e.g., a line chart dataset containing only monotonic trends). To ensure outputs encompass diverse perspectives, it is crucial to identify and rectify these biases. Fairness bias often emerges when models disproportionately rely on their pre-trained knowledge, possibly neglecting the specific visual cues present in a chart. This reliance can lead to outputs that do not accurately reflect the visual data. For instance, if a model is trained on a dataset in which the charts consistently depict a particular group as having a higher diabetes rate compared to others, it is highly probable that when this model encounters a chart presenting contradictory information, it may still generate descriptions that align with its training, favoring interpretations that match the patterns it has learned during pre-training, even if those interpretations are less likely given the new data. Meanwhile, sample bias is evident when datasets over-represent certain types of charts. For example, some models, such as MatCha, exhibit proficiency in interpreting line charts where data points follow a monotonic sequence, but struggle to comprehend line charts characterized by convex or concave configurations, as well as those with fluctuating trends [14]. To address these issues, future work can develop benchmarks to include charts designed to counteract these biases and feature complex configurations, deviating from the typical charts in existing datasets. Moving forward, strategies to mitigate these biases should focus on broadening the diversity of training data and implementing adversarial training techniques, fostering a more equitable and inclusive approach to content generation.
@@ 6 Conclusion
In this survey, we provide a thorough examination of the evolving field of automatic chart understanding, with an emphasis on recent developments such as the transformative impact of large vision-language models. Our review starts with a detailed discussion on the diverse datasets that drive chart understanding research, analyzing their sources, diversity, and the specific tasks they support. We then categorizes the landscape of chart understanding, covering from foundational classification-based and generation-based models to the cutting-edge applications of LVLMs. By offering an in-depth overview of methodologies, critically examining the role of datasets, and outlining future research trajectories, our survey aspires to spur innovation and exploration in the vibrant field of chart understanding, marking a pathway for future breakthroughs in this crucial area of study.

### v4

@@ 4 Modeling Strategies
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, line width=0.9pt, draw=connect-line, rounded corners,align=left, minimum width=2em, s sep=5pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=6em,font=, where level=3font=, where level=4font=, where level=5font=, [ Chart Understanding Approaches, black,rotate=90,anchor=north,edge=pure [ Classification-based Methods (§4.1) , pure-middle, edge=pure, text width=11.7em [ Fixed Output Vocab, pure-middle, edge=pure, text width=10.3em [ IMG+QUES [1], Relation Networks [73], V-MODEQA [45] , pure-leaf, text width=22em, edge=pure ] ] [ Dynamic Encoding, pure-middle, edge=pure, text width=10.3em [ SANDY [1], PreFIL [74], LEAF-Net [43], PlotQA-Pipe [4] , pure-leaf, text width=22em, edge=pure ] ] [ Small Pre-trained Models, pure-middle, edge=pure, text width=10.3em [ STL-CQA [44], TasPas [75], VisionTaPas [3] , pure-leaf, text width=22em, edge=pure ] ] ] [ Generation-based Methods (§4.2), node-middle, edge=node, text width=11.7em [ Without Pre-training, node-middle, edge=node, text width=10.3em [ SciCap-Net [6], FigCap-Net [51], Chart2Text [52] , node-leaf, text width=22em, edge=node ] ] [ Pre-trained, node-middle, edge=node, text width=10.3em [ ChartT5 [76], Donut [26], Pix2Struct [77], MatCha [78],
UniChart [79], FigCaps-HF [56], ChartVE [14] , node-leaf, text width=22em, edge=node ] ] [ Large Vision-language Models, node-middle, edge=node, text width=10.3em [ General-purpose, node-middle, edge=node, text width=5.5em [ LLaVA [17], SPHINX [80], mPLUG-Owl [81],
mPLUG-DocOwl [82], InternVL [83],
GPT-4V [16], Gemini [84], Claude 3 [85] , node-leaf, text width=14.7em, edge=node ] ] [ Chart-specific, node-middle, edge=node, text width=5.5em [ MMCA [48], ChartVLM [58], OneChart [86],
ChartAssistant [42], ChartInstruct [59],
TinyChart [87], ChartLlama [41] , node-leaf, text width=14.7em, edge=node ] ] ] ] [ Tool Augmentation (§4.3), graph-middle, edge=graph, text width=11.7em [ OCR-based, graph-middle, edge=graph, text width=10.3em [ CRAFT [88], Parseq [89], RARE [90], FAN [91], ABINet++ [92],
CRNN [93], Star-Net [94], GRCNN [95], Rosetta [96] ,graph-leaf, text width= 22em, edge=graph ] ] [ Non-OCR-based, graph-middle, edge=graph, text width=10.3em [ DePlot [10], Chart-to-Table [14], DOMINO [97], PromptChart [11],
StructChart [60] ,graph-leaf, text width= 22em, edge=graph ] ] ] ]
@@ 4.1 Classification-based Models
Classification-based methods have mainly tackle the task of chart question answering, leveraging a combination of visual and textual features to understand and interpret different types of charts. This subsection provides an overview of the evolution and key strategies employed in classification-based chart understanding.
@@ 4.2.2 Small Pre-trained Models
Significantly advancing the field, end-to-end visual document understanding models like Donut [26] and Pix2Struct [77] introduced the capability to directly parse textual information from images, removing the need for separate OCR processes. Donut utilizes a Swin Transformer [106] for the image encoding and the decoder of mBART [107] for decoding, focusing on natural image parsing to produce text sequences from images. Pix2Struct, employing a Vision Transformer (ViT) [108] as the image encoder and a Transformer-based decoder. The pre-training objective is screenshot parsing, which learns to predict an HTML-based parse from a screenshot of a webpage, providing a structured training signal that encapsulates both textual and layout information. There are two major distinctions between Donut and Pix2Struct: (1) screenshot parsing may allow the model to learn the structure of the image better since HTML provides stronger signals about the layouts of the input image. (2) Donut is multilingual since it uses mBART’s decoder for decoding texts.
@@
First, do we still need external OCR systems? The advancements in chart understanding and interpretation achieved by UniChart and MatCha over ChartT5 highlight a pivotal shift in the field towards leveraging OCR-free models such as Pix2Struct and Donut. These developments underscore a crucial insight: the integration of advanced OCR-free pre-training objectives can obviate the dependency on traditional OCR systems, which often represent a bottleneck in terms of accuracy and adaptability. UniChart and MatCha exhibit superior performance without needing the direct input of ground truth tables, which ChartT5 relies upon. This clearly shows the increased efficacy and robustness of models developed using the OCR-free paradigm.
@@
TABLE IV: Summary of various pre-trained and instruction fine-tuned vision-language models for chart understanding. Note that MMCA is instruction-tuned from mPLUG-Owl, ChartLlama is instruction-tuned from LLaVA-1.5, and ChartAssistant-S is instruction-tuned from SPHINX. The weights of ChartInstruct’s vision encoder are initialized from UniChart.
@@ 4.2.3 Large Vision-language Models
While the aforementioned pre-trained approaches significantly enhance chart understanding, they require task-specific fine-tuning (see Figure 3), limiting their generalizability. Contrastingly, large foundation models have demonstrated strong zero-shot or few-shot capabilities. For instance, in text generation, LLMs like GPT-3 [40] showcase remarkable generalization to novel tasks when the model size and pre-training data reach certain scales. These models are able to handle new tasks in a zero-shot or few-shot manner, obviating the need for model retraining. Extending this paradigm into image understanding, studies have stacked vision encoders on top of LLMs, including methods such as LLaVA [17] and mPLUG-Owl [81], significantly advancing this field. These larger vision-language models (LVLMs) differ from the smaller vision-language models introduced previously by having substantially greater model sizes and instruction tuning. Unique to these LVLMs is the inclusion of a visual-language projector, $\textsc{Projector}_{\text{vt}}$, given the vision encoders and LLMs’ pre-training scales. Here, the visual representations obtained from the vision encoder are fed to $\textsc{Projector}_{\text{vt}}$ prior to concatenated with text representations.
@@
Finally, is it possible for models trained exclusively on synthetically generated data to match or exceed the performance of those trained on datasets curated from human-generated content? While synthetic data offers the advantage of scalability and control over diversity and complexity, it may lack the nuanced and unpredictable variations present in human-generated charts. Exploring whether synthetic data can truly capture the breadth of human creativity and intricacy in chart design is essential, as is understanding the potential gaps that might arise when such models are applied to real-world data. Recent work [59] has shown that completely removing real-world data from the training set significantly reduces performance. However, adding synthetic training data helps improve data efficiency and reduces the need for large-scale real-world chart data. Future research can focus on studying the effectiveness of synthetic data at varying scale and on developing methods for creating synthetic data that can better mimic real-world complexities.
@@ 5.2 Evaluation
Fairness and Bias. Evaluating fairness and mitigating biases in the generation of models’ content is critical when deploying these tools in user-facing applications. There are two main forms of bias to be aware of: fairness bias, which concerns the equitable representation and treatment of various groups (e.g., race, culture, or gender) [28, Yang_Yu_Fung_Li_Ji_2023], and sample bias, which occurs when the data used to train the model is not representative of the broader population or scenario it’s intended to serve (e.g., a line chart dataset containing only monotonic trends). To ensure outputs encompass diverse perspectives, it is crucial to identify and rectify these biases. Fairness bias often emerges when models disproportionately rely on their pre-trained knowledge, possibly neglecting the specific visual cues present in a chart. This reliance can lead to outputs that do not accurately reflect the visual data. For instance, if a model is trained on a dataset in which the charts consistently depict a particular group as having a higher diabetes rate compared to others, it is highly probable that when this model encounters a chart presenting contradictory information, it may still generate descriptions that align with its training, favoring interpretations that match the patterns it has learned during pre-training, even if those interpretations are less likely given the new data. Meanwhile, sample bias is evident when datasets over-represent certain types of charts. For example, some models, such as MatCha, exhibit proficiency in interpreting line charts where data points follow a monotonic sequence, but struggle to comprehend line charts characterized by convex or concave configurations, as well as those with fluctuating trends [14]. To address these issues, future work can develop benchmarks to include charts designed to counteract these biases and feature complex configurations, deviating from the typical charts in existing datasets. Future strategies should focus on diversifying training data and using adversarial training techniques to mitigate these biases. This will promote a more equitable and inclusive approach to content generation.
@@ 6 Conclusion
In this survey, we provide a thorough examination of the evolving field of automatic chart understanding, with an emphasis on recent developments such as the transformative impact of large vision-language models. Our review starts with a detailed discussion on the diverse datasets that drive chart understanding research, analyzing their sources, diversity, and the specific tasks they support. We then categorizes the landscape of chart understanding, covering from foundational classification-based and generation-based models to the cutting-edge applications of LVLMs. We offer an in-depth overview of methodologies. Additionally, we critically examine the role of datasets and outline future research trajectories. Our survey aims to spur innovation and exploration in the vibrant field of chart understanding. This marks a pathway for future breakthroughs in this crucial area of study.